OpenStudy (anonymous):

solve: 6(x+4)=3(4x-2)

OpenStudy (anonymous):

so you get 6x+24=12x-6

DivineSolar (divinesolar):

Correct and you want to solve that equation?

OpenStudy (anonymous):

6x+30=12x

OpenStudy (anonymous):

Let's solve your equation step-by-step. 6(x+4)=3(4x−2) Step 1: Simplify both sides of the equation. 6(x+4)=3(4x−2) Simplify: (Show steps) 6x+24=12x−6 Step 2: Subtract 12x from both sides. 6x+24−12x=12x−6−12x −6x+24=−6 Step 3: Subtract 24 from both sides. −6x+24−24=−6−24 −6x=−30 Step 4: Divide both sides by -6. −6x −6 = −30 −6 x=5 Answer: x=5

DivineSolar (divinesolar):

@amistre64